# A unique Neutrino Mass Matrix Texture under Exponential Parametrization

### v2

@@ Appendix B The Scalar Potential
Without loss of generality, the following equations are valid from the minimisation conditions of the scalar potential for the chosen vacuum alignments: $\langle H\rangle_{0}=v_{H}(1,1,1)^{T}$, $\langle\Delta\rangle_{0}=v_{\Delta}(v_{1},v_{2},0)^{T}$, $\langle\chi\rangle_{0}=v_{\chi}$, $\langle\psi\rangle_{0}=v_{\psi}$, $\langle\xi\rangle_{0}=v_{\xi}$, $\langle\eta\rangle_{0}=v_{\eta}$ and $\langle\kappa\rangle_{0}=v_{\kappa}$,
$v_{H}(\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+v_{1}^{2}(\lambda_{1}^{H\Delta}+2\lambda_{2}^{H\Delta})+2v_{1}v_{2}(\lambda_{3}^{H\Delta}-\lambda_{4}^{H\Delta})+v_{2}(2\mu+\lambda_{1}^{H\Delta}v_{2}$
$-\lambda_{2}^{H\Delta}v_{2})+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}v_{\kappa}^{2}+\lambda_{1}^{H\xi}v_{\xi}^{2}+\lambda_{1}^{H\chi}v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0,$ (46)
$v_{H}(\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+v_{1}^{2}(\lambda_{1}^{H\Delta}-\lambda_{2}^{H\Delta})+2v_{1}v_{2}(\lambda_{3}^{H\Delta}+\lambda_{4}^{H\Delta})+v_{2}^{2}(\lambda_{1}^{H\Delta}+2\lambda_{2}^{H\Delta})$
$+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}v_{\kappa}^{2}+\lambda_{1}^{H\xi}v_{\xi}^{2}+\lambda_{1}^{H\chi}v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0$ (47)
$v_{H}(\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+(\lambda_{1}^{H\Delta}-\lambda_{2}^{H\Delta})(v_{1}^{2}+v_{2}^{2})+2\mu v_{1}+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}\,v_{\kappa}^{2}+\lambda_{1}^{H\xi}\,v_{\xi}^{2}+$
$\lambda_{1}^{H\chi}\,v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0,$ (48)
$v_{H}^{2}(3\lambda_{1}^{H\Delta}v_{1}+2(\mu+\lambda_{3}^{H\Delta}v_{2}-\lambda_{4}^{H\Delta}v_{2}))+v_{1}(\mu_{\Delta}^{2}+2\lambda_{1}^{\Delta}v_{1}^{2}+2v_{2}^{2}(\lambda_{1}^{\Delta}+2\lambda_{3}^{\Delta}-\lambda_{4}^{\Delta})+\lambda_{1}^{\Delta\eta}v_{\eta}^{2}$
$+\lambda_{1}^{\Delta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\Delta\xi}v_{\xi}^{2}+\lambda_{1}^{\Delta\chi}v_{\chi}^{2}+\lambda_{1}^{\Delta\psi}v_{\psi}^{2})=0,$ (49)
$v_{H}^{2}(2(\mu+v_{1}(\lambda_{3}^{H\Delta}+\lambda_{4}^{H\Delta}))+3\lambda_{1}^{H\Delta}v_{2})+v_{2}(\mu_{\Delta}^{2}+2v_{1}^{2}(\lambda_{1}^{\Delta}+2\lambda_{3}^{\Delta}+\lambda_{4}^{\Delta})+2\lambda_{1}^{\Delta}v_{2}^{2}$
$+\lambda_{1}^{\Delta\eta}v_{\eta}^{2}+\lambda_{1}^{\Delta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\Delta\xi}v_{\xi}^{2}+\lambda_{1}^{\Delta\chi}v_{\chi}^{2}+\lambda_{1}^{\Delta\psi}v_{\psi}^{2})=0,$ (50)
@@
$v_{\chi}(\mu_{\chi}^{2}+3\lambda_{1}^{H\chi}v_{H}^{2}+\lambda_{1}^{\Delta\chi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\chi\eta}v_{\eta}^{2}+\lambda_{1}^{\chi\kappa}v_{\kappa}^{2}+\lambda_{1}^{\chi\xi}v_{\xi}^{2}+2\lambda_{1}^{\chi}v_{\chi}^{2}+\lambda_{1}^{\chi\psi}v_{\psi}^{2})=0,$ (52)
$v_{\psi}(\mu_{\psi}^{2}+3\lambda_{1}^{H\psi}v_{H}^{2}+\lambda_{1}^{\Delta\psi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}{\psi\eta}v_{\eta}^{2}+\lambda_{1}^{\psi\kappa}v_{\kappa}^{2}+\lambda_{1}^{\psi\xi}v_{\xi}^{2}+\lambda_{1}^{\chi\psi}v_{\chi}^{2}+2\lambda_{\psi}v_{\psi}^{2})=0,$ (53)
$v_{\xi}(\mu_{\xi}^{2}+3\lambda_{1}^{H\xi}v_{H}^{2}+\lambda_{1}^{\Delta\xi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\xi\eta}v_{\eta}^{2}+\lambda_{1}^{\xi\kappa}v_{\kappa}^{2}+2\lambda_{\xi}v_{\xi}^{2}+\lambda_{1}^{\chi\xi}v_{\chi}^{2}+\lambda_{1}^{\psi\xi}v_{\psi}^{2})=0,$ (54)
$v_{\eta}(\mu_{\eta}^{2}+2\lambda_{1}^{\Delta\eta}(v_{1}^{2}+v_{2}^{2})+2\lambda^{\eta}v_{\eta}^{2}+\lambda_{1}^{\eta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\xi\eta}v_{\xi}^{2}+\lambda_{1}^{\chi\eta}v_{\chi}^{2}+\lambda_{1}^{\psi\eta}v_{\psi}^{2}+3\lambda_{1}^{H\eta}v_{H}^{2})=0,$ (55)
$v_{\kappa}(\mu_{\kappa}^{2}+2\lambda_{1}^{\Delta\kappa}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\eta\kappa}v_{\eta}^{2}+2\lambda^{\kappa}v_{\kappa}^{2}+\lambda_{1}^{\xi\kappa}v_{\xi}^{2}+\lambda_{1}^{\chi\kappa}v_{\chi}^{2}+\lambda_{1}^{\psi\kappa}v_{\psi}^{2}+3\lambda_{1}^{H\kappa}v_{H}^{2})=0.$ (56)

### v3

@@ Appendix B The Scalar Potential
For simplicity, we redefine $\mu=\frac{1}{\Lambda}\mu_{H\Delta}\xi$. Without loss of generality, the following equations are valid from the minimisation conditions of the scalar potential for the chosen vacuum alignments: $\langle H\rangle_{0}=v_{H}(1,1,1)^{T}$, $\langle\Delta\rangle_{0}=v_{\Delta}(v_{1},v_{2},0)^{T}$, $\langle\chi\rangle_{0}=v_{\chi}$, $\langle\psi\rangle_{0}=v_{\psi}$, $\langle\xi\rangle_{0}=v_{\xi}$, $\langle\eta\rangle_{0}=v_{\eta}$ and $\langle\kappa\rangle_{0}=v_{\kappa}$,
$v_{H}(-\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+v_{1}^{2}(\lambda_{1}^{H\Delta}+2\lambda_{2}^{H\Delta})+2v_{1}v_{2}(\lambda_{3}^{H\Delta}$
$-\lambda_{4}^{H\Delta})+v_{2}(2\mu+\lambda_{1}^{H\Delta}v_{2}-\lambda_{2}^{H\Delta}v_{2})+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}v_{\kappa}^{2}+\lambda_{1}^{H\xi}$
$v_{\xi}^{2}+\lambda_{1}^{H\chi}v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0,$ (10)
$v_{H}(-\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+v_{1}^{2}(\lambda_{1}^{H\Delta}-\lambda_{2}^{H\Delta})+2v_{1}v_{2}(\lambda_{3}^{H\Delta}$
$+\lambda_{4}^{H\Delta})+v_{2}^{2}(\lambda_{1}^{H\Delta}+2\lambda_{2}^{H\Delta})+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}v_{\kappa}^{2}+\lambda_{1}^{H\xi}v_{\xi}^{2}+$
$\lambda_{1}^{H\chi}v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0$ (11)
$v_{H}(-\mu_{H}^{2}+v_{H}^{2}(6\lambda_{1}^{H}+8\lambda_{3}^{H})+(\lambda_{1}^{H\Delta}-\lambda_{2}^{H\Delta})(v_{1}^{2}+v_{2}^{2})+2\mu v_{1}$
$+\lambda_{1}^{H\eta}v_{\eta}^{2}+\lambda_{1}^{H\kappa}\,v_{\kappa}^{2}+\lambda_{1}^{H\xi}\,v_{\xi}^{2}+\lambda_{1}^{H\chi}\,v_{\chi}^{2}+\lambda_{1}^{H\psi}v_{\psi}^{2})=0,$ (12)
$v_{H}^{2}(3\lambda_{1}^{H\Delta}v_{1}+2(\mu+\lambda_{3}^{H\Delta}v_{2}-\lambda_{4}^{H\Delta}v_{2}))+v_{1}(-\mu_{\Delta}^{2}+2(\lambda_{1}^{\Delta}+\lambda_{2}^{\Delta})v_{1}^{2}$
$+2v_{2}^{2}(\lambda_{1}^{\Delta}-\lambda_{2}^{\Delta}+2\lambda_{3}^{\Delta}-\lambda_{4}^{\Delta})+\lambda_{1}^{\Delta\eta}v_{\eta}^{2}+\lambda_{1}^{\Delta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\Delta\xi}v_{\xi}^{2}+\lambda_{1}^{\Delta\chi}$
$v_{\chi}^{2}+\lambda_{1}^{\Delta\psi}v_{\psi}^{2})=0,$ (13)
$v_{H}^{2}(2(\mu+v_{1}(\lambda_{3}^{H\Delta}+\lambda_{4}^{H\Delta}))+3\lambda_{1}^{H\Delta}v_{2})+v_{2}(-\mu_{\Delta}^{2}+2v_{1}^{2}(\lambda_{1}^{\Delta}-\lambda_{2}^{\Delta}+$
$2\lambda_{3}^{\Delta}+\lambda_{4}^{\Delta})+2(\lambda_{1}^{\Delta}+\lambda_{2}^{\Delta})v_{2}^{2}+\lambda_{1}^{\Delta\eta}v_{\eta}^{2}+\lambda_{1}^{\Delta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\Delta\xi}v_{\xi}^{2}+\lambda_{1}^{\Delta\chi}$
$v_{\chi}^{2}+\lambda_{1}^{\Delta\psi}v_{\psi}^{2})=0,$ (14)
@@
$v_{\chi}(-\mu_{\chi}^{2}+3\lambda_{1}^{H\chi}v_{H}^{2}+\lambda_{1}^{\Delta\chi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\chi\eta}v_{\eta}^{2}+\lambda_{1}^{\chi\kappa}v_{\kappa}^{2}+\lambda_{1}^{\chi\xi}v_{\xi}^{2}$
$+2\lambda_{1}^{\chi}v_{\chi}^{2}+\lambda_{1}^{\chi\psi}v_{\psi}^{2})=0,$ (16)
$v_{\psi}(-\mu_{\psi}^{2}+3\lambda_{1}^{H\psi}v_{H}^{2}+\lambda_{1}^{\Delta\psi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}{\psi\eta}v_{\eta}^{2}+\lambda_{1}^{\psi\kappa}v_{\kappa}^{2}+\lambda_{1}^{\psi\xi}v_{\xi}^{2}$
$+\lambda_{1}^{\chi\psi}v_{\chi}^{2}+2\lambda_{\psi}v_{\psi}^{2})=0,$ (17)
$v_{\xi}(-\mu_{\xi}^{2}+3\lambda_{1}^{H\xi}v_{H}^{2}+\lambda_{1}^{\Delta\xi}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\xi\eta}v_{\eta}^{2}+\lambda_{1}^{\xi\kappa}v_{\kappa}^{2}+2\lambda_{\xi}v_{\xi}^{2}+$
$\lambda_{1}^{\chi\xi}v_{\chi}^{2}+\lambda_{1}^{\psi\xi}v_{\psi}^{2})=0,$ (18)
$v_{\eta}(-\mu_{\eta}^{2}+2\lambda_{1}^{\Delta\eta}(v_{1}^{2}+v_{2}^{2})+2\lambda^{\eta}v_{\eta}^{2}+\lambda_{1}^{\eta\kappa}v_{\kappa}^{2}+\lambda_{1}^{\xi\eta}v_{\xi}^{2}+\lambda_{1}^{\chi\eta}v_{\chi}^{2}+$
$\lambda_{1}^{\psi\eta}v_{\psi}^{2}+3\lambda_{1}^{H\eta}v_{H}^{2})=0,$ (19)
$v_{\kappa}(-\mu_{\kappa}^{2}+2\lambda_{1}^{\Delta\kappa}(v_{1}^{2}+v_{2}^{2})+\lambda_{1}^{\eta\kappa}v_{\eta}^{2}+2\lambda^{\kappa}v_{\kappa}^{2}+\lambda_{1}^{\xi\kappa}v_{\xi}^{2}+\lambda_{1}^{\chi\kappa}v_{\chi}^{2}+$
$\lambda_{1}^{\psi\kappa}v_{\psi}^{2}+3\lambda_{1}^{H\kappa}v_{H}^{2})=0.$ (20)